OpenStudy (yinshy):

What is the 12th term in the addition pattern with formula 6n – 7? A. 59 B. 65 C. 71 D. 72

OpenStudy (yinshy):

@ijlal

OpenStudy (isaiah.feynman):

Let f(n) = 6n - 7 Then first term is f(1) = 6(1) - 7 Second term is f(2) = 6(2) - 7 Third term is f(3) = 6(3) - 7 See the pattern now?

OpenStudy (yinshy):

A?

OpenStudy (ijlal):

it is how Isasiah stated

OpenStudy (ijlal):